OpenStudy (anonymous):

can someone help me write the equation for a circle with a diameter that has endpoints at (–4, –1) and (–8, –9).

OpenStudy (anonymous):

equation of a circle is equal to : \[(x-a)^2 + (y-b)^2 = r^2\] where r = radius and (a,b) are coordinates of the centre of the circle

OpenStudy (anonymous):

here your diameter's end points are (-4,-1) and (-8,-9) meaning :  where c is the centre of the circle, which is also the midpoint of these 2 points